The Fibonacci sequence follows the pattern 1,1,2,3,5,8,13,21,34,55,....which of the following numbers is part of the fibonacci sequence?
F.233
G 378
H 612
J 988

Answer by robertb(5830)

F. 233
The term after 55 is 55+34 = 89, the next is
55 + 89 = 144, the next is
89 + 144 = 233, next is
144 + 233 = 377, next is
233 + 377 = 610, next is
377 + 610 = 987.